# Analyzing Similarity Metrics for Data Selection for Language Model Pretraining

### v1

@@ 1 Introduction
The recent success of general purpose language models (Brown, 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024). Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning. And so, researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from larger corpora of examples scraped from various sources.
@@ 4.1 Experiment Details
To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500000 examples) and extract the top 64 dimensions upon which to project. The projected embeddings are then normalized to have a unit L2 norm. We also explore using random projections (Bingham & Mannila, 2001) as a different dimensionality reduction strategy, finding that random projections do not perform as well as PCA (see Section 4.5 for ablation studies).
@@
For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3. We use a selection budget of 20B tokens, which is approximately 20% of sequences from the Pile. This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
We pretrain with a learning rate of 0.001 with a linear decay and a batch size of 1024. For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens. For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown (2020). This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks. The full details about our evaluation sets are deferred to Section B.1. For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 20B randomly selected tokens.
@@ 4.3 Cluster Purity Results
Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 20B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective. Bolding and underlining denote the best and second-best performing methods on each task, respectively. Results are averaged over 3 pretraining runs. We find that embeddings extracted from small version of the downstream language model and Gecko perform the best for pretraining data curation.
@@ B.3 Additional Hyperparameter Details
For our RAC clustering for pretraining experiments at a 20B token budget, we use a value of $\epsilon$ as the particular diameter of clusters:
@@ B.4 Dimensionality Reduction Details
For running our dimensionality reduction via PCA, we compute the means and components on which to project on a subset of the data ($\sim$500000 points). We first standardize the embeddings to have a mean of 0 and variance of 1 before running PCA. As previously mentioned, after projecting onto the desired number of principal components, we perform L2 normalization.

### v2

@@ 1 Introduction
The recent success of general purpose language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024). Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning. And so, researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from larger corpora of examples scraped from various sources.
@@ 4.1 Experiment Details
To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500,000 examples) and extract the top 64 dimensions upon which to project. The projected embeddings are then normalized to have a unit L2 norm. We also explore using random projections (Bingham & Mannila, 2001) as a different dimensionality reduction strategy, finding that random projections do not perform as well as PCA (see Section 4.5 for ablation studies).
@@
For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3. We use a selection budget of 200B tokens, which is approximately 20% of sequences from the Pile. This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
We pretrain with a learning rate of 0.001 with a linear decay and a batch size of 1024. For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens. For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown et al. (2020). This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks. The full details about our evaluation sets are deferred to Section B.1. For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 200B randomly selected tokens.
@@ 4.3 Cluster Purity Results
Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 200B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective. Bolding and underlining denote the best and second-best performing methods on each task, respectively. Results are averaged over 3 pretraining runs. We find that embeddings extracted from small version of the downstream language model and Gecko perform the best for pretraining data curation.
@@ B.3 Additional Hyperparameter Details
For performing RAC clustering for our pretraining experiments, we use a value of $\epsilon$ as the particular diameter of clusters:
@@ B.4 Dimensionality Reduction Details
For running our dimensionality reduction via PCA, we compute the means and components on which to project on a subset of the data ($\sim$500,000 points). We first standardize the embeddings to have a mean of 0 and variance of 1 before running PCA. As previously mentioned, after projecting onto the desired number of principal components, we perform L2 normalization.